# Integration of a high-fidelity model of quantum sensors with a map-matching filter for quantum-enhanced navigation

### v1

@@ III.2 Quantum sensor specifications
In all simulations, we have considered a m-baseline gravity gradiometer, with atoms in each cloud and a total interrogation time . Other sensor parameters including cloud temperature and diameter , laser power and beam diameter , have been varied broadly to simulate sensors from lower- to higher- technological readiness. In the following, three typical scenarios will be referred to to benchmark the discussion of the results: (i) a ”moderate” scenario, representative of current sensor specifications: , , , , (ii) an ”enhanced” scenario, representative of potential future sensor specifications: , , , ; (iii) a ”enhanced + mitigated” scenario, where we have assumed that in addition to the sensor specifications of the ”enhanced” scenario, additional mitigation techniques, such as a gimbal platform or rotation compensation techniques, would permit to reduce sensor tilts and rotation rates by an arbitrary and tunable numerical factor.
@@
Each gravity gradient measurement is obtained by fitting an ellipse made of 15 data points 222About 15-25 shots per ellipse has been shown to be an optimum - large enough to avoid standard errors and small enough to avoid unnecessarily increasing measurement time Stray et al. (2022). These 15 shots are successively obtained by running our model while chirping the laser frequency. As the sensor is moving along the trajectory from one shot to the following one Phillips et al. (2022), shot repetition rate is an important parameter. To allow investigation of fast repetition rates, without being limited by the actual duration of the interferometric sequence (ms interrogation time, without adding MOT loading times), we have neglected variations of the sensor kinematic parameters (latitude/longitude, velocity, acceleration, attitude and rotation rates) within each shot, allowing us to simulate shots as if they happened instantaneously. Therefore, our simulations at fast repetition rates do not imply that we have reduced the interrogation time, which remains ms all throughout. This situation could be representative of using advanced techniques such as interleaving measurements using several sensors or improved detection methods Sugarbaker et al. (2013). In the following, unless stated otherwise, the repetition rate between shots is 1.5Hz for trajectory 1 (hence a 0.1Hz gravity gradient measurement sample rate) and 15Hz for trajectories 2 and 3 (1Hz gravity gradient measurement sample rate).
@@ III.3 Quantum measurements
In the ”enhanced” specifications scenario (a), the measurements display significant errors all throughout the trajectory, that decrease gradually as the ship goes through calmer sea states. It is is only in sea state 1 (s), that the gravity gradient signal is not completely written out by noise [see inset]. Conducting additional simulations while isolating out different contributions of the ship motion has shown that these errors were imputable to both instantaneous rotation rates and rotation change rates.
Both significantly contribute to the interferometer phase shift, leading to a systematic error in the estimated gravity gradient that, due to the rapid change in these quantities shot-to-shot, appears as noise.
@@
In turn, static tilts and ship transverse acceleration were found to have a limited effect, although in sea state 4 we do start to see effects of pulse infidelities arising from the atoms shifting off-centered in the beam as a result of tilt and centripetal force. Those effects would be more pronounced if assuming the sensor was positioned further away from the platform centre of rotation [see discussion in Appendix B].
@@
In sea state 1, where the ”enhanced” scenario (a) provides good measurements, we have also investigated the effect of reducing sensor technological readiness. The results of the ”moderate” specification scenario, shown in Fig. 4 (c), evidence an increase of noise compared to the ”enhanced” scenario as a result of cross-coupling effects arising from atomic cloud inhomogeneities (hotter, larger cloud). Interestingly though, this additional noise arising from inter-atomic cross-couplings fundamentally differs from the previously described inertial errors in the sense that it manifests itself as Gaussian noise.
Finally, Fig. 4 (d) shows simulations of the ”enhanced” scenario with a higher repetition rate of 10Hz per shot (hence a 0.67Hz gradiometry sample rate). We observe that measurements are slightly noisier, which here does not arise via a reduced interrogation time, but rather via the fact that high data-rate measurements are more sensitive to high-frequency rotational noise from waves. Indeed, when using a lower data-rate, changes in inertial quantities between consecutive ellipse points result in these points sitting on different ellipses - producing a noisier ellipse, yet whose fitting smooths out these effects.
@@
Figure 5 displays the simulated gravity gradient measurements along trajectories 2 and 3. Even in the ”enhanced” scenario (top plots), the gravity gradient measurements are severely deteriorated. For trajectory 2, we find that the significant systematic shift occurring for s is due to variations of the roll angle, which create a rotational phase shift in the interferometer that largely exceeds the gravity-gradient induced phase shift; the measurement curve closely follows the variations of the roll angle, and the sensor actually measures a rotation rate instead of a gravity gradient. For s, sensor tilts, accelerations and rotations are all responsible for the observed noise, each resulting in significant errors even when taken out individually. Similar observations are made for trajectory 3, where large systematic shifts are imputed to variations of the roll and pitch angles, and, for s, to large sensor tilts. For both trajectories, the ”enhanced + mitigated” scenario, which we have simulated here with a 100-fold reduction in roll and pitch angles, shows a highly improved performance, with a 1 gradiometry residual noise below over most of the trajectories.
@@ IV Integration with a map-matching navigation filter
In this section, we demonstrate the integration of gradiometry digital measurements within an error-state Integrated Navigation filter. Gradiometry measurements are first processed through a map-matching algorithm to generate periodic horizontal position estimates. This estimate is then passed to the navigation filter, which provides position, velocity and attitude corrections to the inertial navigation solution derived from the IMU. The IMU errors assumed here are: for trajectory 1 (ship), an accelerometer bias of g, scale factor error of 25ppm and misalignment of rad, and a gyroscope bias of /hr, scale factor error of 5ppm and misalignment of rad; for trajectories 2 and 3 (flight), an accelerometer bias of 3mg, scale factor error of 1000ppm and misalignment of 0.5mrad, and a gyroscope bias of /hr, scale factor error of 1000ppm and misalignment of 0.5mrad. All simulations start with 30s of Transfer Alignment (TFA) which prime the navigation solution and during which gradiometric measurements are not utilized to initialise the system position, velocity and attitude states. A TFA initialisation error of 100m North and 100m East is assumed in all cases.
@@
As this work focuses on demonstrating successful integration of quantum and map-matching capabilities rather than maximizing ultimate navigation performance, we have prioritized generic tunings that ensure covariance consistency across all measurement capabilities over platform- and sensor-specific optimization.
@@
Similar conclusions were obtained for sea states 3 and 4, as well as for flight trajectories 2 and 3, where, due to large systematic errors, only the ”enhanced + mitigated” scenario could be used to exercise the map-matching algorithm. However, due to the large errors present, we have found that more severe mitigation was required.
Fig. 7 shows map-matching results for trajectories 2 and 3, assuming a 100-fold reduction in tilts/rotations. With position errors remaining bounded over time, it shows that such ”enhanced + mitigated” gradiometric measurements at 1Hz are sufficient to allow efficient position fixing even in high-dynamic platforms.
@@ V Discussion
These criteria are found to be generally consistent with what is observed in Sec. III.3 and with the levels of successful mitigation evidenced for the various platforms considered in this work.
@@
However, especially on future air and compact marine platforms where the use of a gimbal sensor may not be practical, a strap-down sensor might be required. Different areas to progress on the mitigation of inertial errors in strap-down sensors may be pursued in the future, including: (i) Developing resilient hardware, such as driving the interferometer using composite Saywell et al. (2018, 2020, 2023) or polychromatic Lellouch et al. (2022) pulses, which have shown promises to enhance the sensor’s resilience to external field and atomic cloud inhomogeneities. Such pulses could be tailored to mitigate the effects induced by tilts, accelerations and rotations using robust and optimal quantum control techniques Glaser et al. (2015). High-bandwidth sensors Rakholia et al. (2014); Lee et al. (2022); Adams et al. (2023) operating at small interrogation times could also help reduce the impact of centripetal and Coriolis effects. (ii) Post-processing techniques to compensate systematic errors out of the gradiometry measurements using additional information such as IMU data. While hybridization between a classical IMU and an atom interferometer has been demonstrated in other navigation contexts Tennstedt et al. (2023), a concept for gravity retrieval in an atom interferometer hybridized with an electrostatic accelerometer has been proposed in Zahzam et al. (2022) and separation between rotation and acceleration signals in an atom interferometer hybridized with a classical sensor has been demonstrated in de Castanet et al. (2024). To be effective in regimes where cross-couplings are significant, such compensation methods would require as accurate an interferometer model as the one presented here. They would also need overcoming potential issues related to non-linearity in fringe reconstruction or underpinning sensor hardware.
@@
While we have discussed errors induced by dynamic conditions of operations, other external factors, such as magnetic fields, may induce similar errors. Effective suppression of DC magnetic effects has been demonstrated in field gravity gradiometers via the use of passive shielding, magnetic coiling and the implementation of techniques such as interleaving measurements Stray (2021). Their implementation on moving platforms, especially flight applications, is mostly dependent upon meeting SWAP constraints. Interestingly, as map-matching algorithms are highly tolerant of Gaussian noise, reducing quantum sensor SWAP could still bring benefits for navigation despite coming with increased noise from inter-atom cross-couplings.
@@
Our integration of a quantum sensor model with a navigation filter has also demonstrated how gradiometry measurement errors impact navigation filter requirements. Rather than seeking to suppress/compensate these errors, an alternative approach would be to make map-matching algorithms more tolerant to them, such as by removing the underlying Gaussian noise assumption. More generally, while filter tuning and quantum sensing error mitigation are traditionally considered separately, our results have highlighted how the optimization of the navigation and the quantum sensor should be considered together. To make the filter as responsive as possible to measurements without risking to become covariant inconsistent, optimal filter tuning should take into account the uncertainty in quantum measurements, which strongly depends on sensor specifications and platform dynamics. Meanwhile, finding the optimal resampling time that most efficiently combats covariance growth is expected to be heavily influenced by the measurement uncertainty, the map roughness and the platform velocity (which influence the chances for a unique position fix), but also by the gradiometer measurement rate; optimally tuning the latter involves a non-trivial trade-off between ensuring enough measurements have been processed to achieve convergence, and noisier quantum measurements/risk of duplicated measurements. In that respect, integrating a quantum sensor model with a navigation filter represents a first critical step on the way to finding the optimum measurement configuration and filter tuning that would be unique to the platform kinematics, map roughness, quantum sensor specifications, and depend on end-user performance requirements.
@@ Acknowledgements:
Appendix A Two-photon Raman transitions
In a stimulated Raman transition, two hyperfine ground-states and are coupled by two laser beams whose frequencies are chosen to be close to the resonant frequencies between those states and a common excited state , ; the detunings and [see Fig. 8] are referred to as the single- and two-photon detuning. If , one can show that state remains mostly unpopulated while the dynamics of the two-level system [, ] is governed by the 2x2 matrix
@@ Appendix A Two-photon Raman transitions
where the detuning is defined below and the effective Rabi frequency is given by Berman (1997); Weiss et al. (1994) with the magnetic quantum number and the Rabi frequency associated with laser j; the latter is given by with the local intensity of laser j at the instantaneous atomic position , the speed of light, the vacuum dielectric permittivity and the transition dipole matrix element;
where is the differential AC Stark-shift and is the Doppler frequency shift due to the atom’s motion, with the transferred momentum and the instantaneous atomic velocity. In this work, we will assume that the laser frequencies are chosen so that an atom at rest is on resonance, , and that the laser intensity ratio between the two beams is set so that ;
, assuming perferctly flat laser equiphases.

### v2

@@ III.2 Quantum sensor specifications
In all the simulations, we have assumed a m-baseline gravity gradiometer, with atoms in each cloud, a total interrogation time and a ms time-of-flight before the first pulse. The situation we consider is that of a well-optimised gravity gradiometer in the laboratory that is put into a navigation environment. In particular, the two counter-propagating lasers are assumed to be generated independently to suppress the effects of parasitic Raman transitions Carraz et al. (2012), and their intensity ratio set to cancel out the AC Stark shift. We assume that effective shielding/ magnetic coiling allows to neglect the effects of magnetic fields, and we assume perfectly flat wavefronts. Other sensor parameters including cloud temperature and diameter , laser power and beam diameter , have been varied broadly to simulate sensors from lower- to higher- technological readiness. In the following, three typical scenarios will be referred to to benchmark the discussion of the results: (i) a ”moderate” scenario, representative of current sensor hardware: , , , , (ii) an ”enhanced” scenario, with specifications commensurate with current high-performance cold-atom experiments, that could be representative of sensor hardware in the close future: , , , ; physically, these parameters correspond to a situation with negligible inter-atom dephasings, where a narrow spread of atomic velocities reduces rotation-induced contrast loss arising from the imperfect overlap between atomic wavepackets, as well-described by the approximate formula Barrett et al. (2016b); (iii) a ”enhanced + mitigated” scenario, where we have assumed that in addition to the sensor specifications of the ”enhanced” scenario, additional mitigation techniques, such as a gimbal platform or rotation compensation techniques, would permit to reduce sensor tilts and rotation rates by an arbitrary and tunable numerical factor.
@@
Each gravity gradient measurement is obtained by fitting an ellipse made of 15 data points–each corresponding to one measurement cycle of the atom interferometer 111About 15-25 shots per ellipse has been shown to be an optimum - large enough to avoid standard errors and small enough to avoid unnecessarily increasing measurement time Stray et al. (2022). These 15 shots are successively obtained by running our model with different values of the chirp rate. The laser wavevector is kept unchanged, as we are not looking into effects that can be suppressed by the k-reversal protocol Gouët et al. (2008). As the sensor is moving along the trajectory from one shot to the following one Phillips et al. (2022), shot repetition rate is an important parameter. To allow investigation of fast interferometric repetition rates, without being limited by the actual duration of the interferometric sequence (ms interrogation time, without adding MOT loading times), we have neglected variations of the sensor kinematic parameters (latitude/longitude, velocity, acceleration, attitude and rotation rates) within each shot of the interferometer, allowing us to simulate each ellipse point as if it happened instantaneously. This technically allows us to repeat the simulation at interferometric repetition rates ms even though the interrogation time remains fixed to ms in all our simulations. This situation could be representative of using advanced techniques such as interleaving measurements using several sensors or improved detection methods Sugarbaker et al. (2013). In the following, unless stated otherwise, the interferometer repetition rate is 1.5Hz for trajectory 1 (hence a 0.1Hz gravity gradient measurement sample rate) and 15Hz for trajectories 2 and 3 (1Hz gravity gradient measurement sample rate). On moving platforms, gravity gradient measurements can be impeded by a reduced fringe contrast. In this work, all extracted data points correspond to ellipses with a semi-major axis of above 0.4.
@@ III.3 Quantum measurements
In the ”enhanced” specifications scenario (a), the measurements display significant errors all throughout the trajectory, that decrease gradually as the ship goes through calmer sea states. It is is only in sea state 1 (s), that the gravity gradient signal is not completely written out by noise [see inset]. Conducting additional simulations while isolating out different contributions of the ship motion has shown that these errors were imputable to both instantaneous rotation rates and rotational accelerations.
Both significantly contribute to the interferometer phase shift, leading to a systematic error in the estimated gravity gradient that, due to the rapid change in these quantities between consecutive gravity gradient measurements, appears as noise.
@@
In turn, static tilts and ship transverse acceleration were found to have a limited effect, although in sea state 4 we do start to see effects of pulse infidelities arising from the atoms shifting off-centered in the beam as a result of tilt and centripetal force. Those effects would be more pronounced if assuming the sensor was positioned further away from the platform centre of rotation [see discussion in Appendix B.2].
Finally, in rough sea states, high noise in the elliptical signal and large phase shifts approaching 0 or give rise to systematic ellipse fitting errors that are a contributing factor to the observed non-Gaussian noise. It is possible to circumvent this via active control of the ellipse phase, which is routinely performed in static differential interferometers but not yet demonstrated in the public domain at sea, so in order not to underestimate the implications of these effects, we have not employed such techniques here.
@@
We have also investigated the effect of reducing sensor technological readiness through considering the ”moderate” specification scenario. As shown in Fig. 5 (c) in the case of calm sea states, it evidences an increase of noise compared to the ”enhanced” scenario as a result of cross-coupling effects arising from atomic cloud inhomogeneities (hotter, larger cloud). Interestingly though, this additional noise arising from inter-atomic cross-couplings fundamentally differs from the previously described inertial errors in the sense that it manifests itself as Gaussian noise. In rougher sea states 3 and 4, we have found that with “moderate” sensor specifications, the contrast reduces to a level which limits the accuracy of current ellipse extraction techniques.
Finally, Fig. 5 (d) shows simulations of the ”enhanced” scenario with a higher interferometer repetition rate of 10Hz per shot (hence a 0.67Hz gradiometry sample rate). We observe that measurements are slightly noisier, which here does not arise via a reduced interrogation time, but rather via the fact that high data-rate measurements are more sensitive to high-frequency rotational noise from waves. Indeed, when using a lower data-rate, changes in inertial quantities between consecutive ellipse points result in these points sitting on different ellipses - producing a noisier ellipse, yet whose fitting smooths out these effects.
@@
Figure 6 displays the simulated gravity gradient measurements along trajectories 2 and 3. Even in the ”enhanced” scenario (top plots), the gravity gradient measurements are severely deteriorated. For trajectory 2, we find that the significant systematic shift occurring for s is due to variations of the roll angle, which create a rotational phase shift [see Appendix B] in the interferometer that largely exceeds the gravity-gradient induced phase shift; we find that the measurement curve on Fig. 6 (top left) closely follows the variations of the roll angle, as observed through on Fig. 4, so that the sensor actually measures a rotation rate instead of a gravity gradient. For s, sensor tilts, accelerations and rotations are all responsible for the observed noise, each resulting in significant errors even when taken out individually. Similar observations are made for trajectory 3, where large systematic shifts are imputed to variations of the pitch angle primarily and, for s, to large sensor tilts in the roll direction. For both trajectories, the ”enhanced + mitigated” scenario, which we have simulated here with a 100-fold reduction in roll and pitch angles, shows a highly improved performance, with a 1 gradiometry residual noise below over most of the trajectories. We note that due to the level of errors observed in the “enhanced” scenario, the “moderate” scenario has not been simulated in flight trajectories.
@@ IV Integration with a map-matching navigation filter
In this section, we demonstrate the integration of gradiometry digital measurements within an error-state Integrated Navigation filter. For the sake of simplicity here, we will only be interested in 2D position fixing. Gradiometry measurements are first processed through a map-matching algorithm to generate periodic horizontal position estimates. This estimate is then passed to the navigation filter, which provides position, velocity and attitude corrections to the inertial navigation solution derived from the Inertial Measurement Unit (IMU). In this work, we are considering tactical/platform grade IMUs, which are among the most commonly used grades for high-capability navigation, such as for GPS holdover, across a broad range of navigation use cases –although not state-of-the-art (e.g. strategic or navigational grade). Therefore, we have assumed the following IMU errors: for trajectory 1 (ship - platform grade), an accelerometer bias of g, scale factor error of 25ppm and misalignment of rad, and a gyroscope bias of /hr, scale factor error of 5ppm and misalignment of rad; for trajectories 2 and 3 (flight - tactical grade), an accelerometer bias of 3mg, scale factor error of 1000ppm and misalignment of 0.5mrad, and a gyroscope bias of /hr, scale factor error of 1000ppm and misalignment of 0.5mrad. All simulations start with 30s of Transfer Alignment (TFA) which prime the navigation solution and during which gradiometric measurements are not utilized to initialise the system position, velocity and attitude states. A TFA initialisation error of 100m North and 100m East is assumed in all cases.
@@
As this work focuses on demonstrating successful integration of quantum and map-matching capabilities, and assessing key filter input requirements, rather than maximizing ultimate navigation performance, we have prioritized generic tunings that ensure covariance consistency across all measurement capabilities over platform- and sensor-specific optimization.
@@
Similar conclusions were obtained for sea states 3 and 4, as well as for flight trajectories 2 and 3, where, due to large systematic errors, only the ”enhanced + mitigated” scenario could be used to exercise the map-matching algorithm. However, due to the large errors present, we have found that more severe mitigation was required. Fig. 8 shows map-matching results for trajectories 2 and 3, assuming a 100-fold reduction in tilts/rotations. With position errors remaining bounded over time, it shows that such ”enhanced + mitigated” gradiometric measurements at 1Hz are sufficient to allow efficient position fixing even in high-dynamic platforms.
@@ V Discussion
These criteria are found to be generally consistent with what is observed in Sec. III.3 and with the levels of successful mitigation evidenced for the various platforms considered in this work given the rotation rates and rotational accelerations shown on Fig. 4.
@@
However, especially on future air and compact marine platforms where the use of a gimbal sensor may not be practical, a strap-down sensor might be required. Different areas to progress on the mitigation of inertial errors in strap-down sensors may be pursued in the future, including: (i) Developing resilient hardware, such as driving the interferometer using composite Saywell et al. (2018, 2020, 2023) or polychromatic Lellouch et al. (2022) pulses, which have shown promises to enhance the sensor’s resilience to external field and atomic cloud inhomogeneities. Such pulses could be tailored to mitigate the effects induced by tilts, accelerations and rotations using robust and optimal quantum control techniques Glaser et al. (2015). High-bandwidth sensors Rakholia et al. (2014); Lee et al. (2022); Adams et al. (2023) operating at small interrogation times could also help reduce the impact of centripetal and Coriolis effects. (ii) Post-processing techniques to compensate systematic errors out of the gradiometry measurements using additional information such as IMU data. While hybridization between a classical IMU and an atom interferometer has been demonstrated in other navigation contexts Tennstedt et al. (2023), a concept for gravity retrieval in an atom interferometer hybridized with an electrostatic accelerometer has been proposed in Zahzam et al. (2022) and separation between rotation and acceleration signals in an atom interferometer hybridized with a classical sensor has been demonstrated in d’Armagnac de Castanet et al. (2024). To be effective in regimes where cross-couplings are significant, such compensation methods would require as accurate an interferometer model as the one presented here. They would also need overcoming potential issues related to non-linearity in fringe reconstruction or underpinning sensor hardware.
@@
While we have focused on errors induced by dynamic conditions of operations, other external factors, such as magnetic fields, may induce similar non-Gaussian errors. Effective suppression of DC magnetic effects has been demonstrated in field gravity gradiometers via the use of passive and active shielding, magnetic coiling and the implementation of techniques such as interleaving measurements Stray (2021). Their implementation on moving platforms, especially flight applications, is mostly dependent upon meeting SWAP constraints. Interestingly, as map-matching algorithms are highly tolerant of Gaussian noise, reducing quantum sensor SWAP could still bring benefits for navigation despite coming with increased noise from inter-atom cross-couplings.
More generally, couplings between the dynamic environment and common systematic effects in atom interferometers, such as light-shifts, Zeeman shifts or wavefront aberration, may induce additional errors that would need further scrutiny. As we have considered a well-optimised gravity gradiometer, with these effects efficiently mitigated in the lab, which is then put into a navigation environment, we would expect these systematic effects to be broadly canceled out. However, possible residual contributions arising from the interplay with the dynamic environment (e.g. dynamical light-shifts, non-inertial atomic motion in aberrated wavefronts…) could give rise to non-trivial errors. While these would not affect the general conclusions of this work, such as the distinction between Gaussian and non-Gaussian errors, they may impact the exact performance, and a detailed analysis of these couplings may constitute the object of a follow-up study.
@@
Our integration of a quantum sensor model with a navigation filter has also demonstrated how gradiometry measurement errors impact navigation filter requirements. Rather than seeking to suppress/compensate these errors, an alternative approach would be to make map-matching algorithms more tolerant to them, such as by removing the underlying Gaussian noise assumption. More generally, while filter tuning and quantum sensing error mitigation are traditionally considered separately, our results have highlighted how the optimization of the navigation and the quantum sensor should be considered together. To make the filter as responsive as possible to measurements without risking to become covariant inconsistent, optimal filter tuning should take into account the uncertainty in quantum measurements, which strongly depends on sensor specifications and platform dynamics. Meanwhile, finding the optimal resampling time that most efficiently combats covariance growth is expected to be heavily influenced by the measurement uncertainty, the map roughness and the platform velocity (which influence the chances for a unique position fix), but also by the gradiometer measurement rate; optimally tuning the latter involves a non-trivial trade-off between ensuring enough measurements have been processed to achieve convergence, and noisier quantum measurements/risk of duplicated measurements.
In that respect, integrating a quantum sensor model with a navigation filter represents a first critical step on the way to finding the optimum measurement configuration and filter tuning that would be unique to the platform kinematics, map roughness, quantum sensor specifications, and depend on end-user performance requirements.
@@ Acknowledgements:
Appendix A Two-photon Raman transitions and application to the 87Rb atom interferometer considered in this work
In a stimulated Raman transition, two hyperfine ground-states and are coupled by two laser beams whose frequencies are chosen to be close to the resonant frequencies between those states and a common excited state , ; the detunings and [see Fig. 9] are referred to as the single- and two-photon detuning. Defining the Rabi frequency associated with laser j, with the intensity of laser j, the speed of light, the vacuum dielectric permittivity and the transition dipole matrix element, one can show that if , state remains mostly unpopulated and the dynamics of the system reduces to that of a two-level system [, ] governed by the 2x2 matrix
@@ Appendix A Two-photon Raman transitions and application to the 87Rb atom interferometer considered in this work
where the detuning is defined below and the effective Rabi frequency is given by Berman (1997). In this work, the single-photon detuning is set to 2.1GHz, and the Rabi frequencies depend on the instantaneous atomic position through the relation with the local intensity of laser j at the instantaneous atomic position .
where is the differential AC Stark-shift and is the Doppler frequency shift due to the atom’s motion, with the effective wavevector and the instantaneous atomic velocity.
is the differential phase between the two Raman beams.
In this work, we assume the two laser beams to come from separate, counter-propagating monochromatic lasers, whose frequencies are chosen so that an atom at rest is on resonance, .
A linearly increasing chirp is further applied to as the cloud free-falls to compensate for the gravity-induced Doppler shift, while varying the chirp rate is used to scan the interferometric fringe.
We assume both lasers to have the same Gaussian intensity distribution, and the intensity ratio between the two beams to be set so that . For simplicity, we assume perferctly flat laser equiphases, .
In practice, a real 87Rb atom features a more complex atomic state manifold. Due to the hyperfine splitting of the excited state, the stimulated Raman transition is no longer strictly a three-level process, as magnetic sub-levels create multiple possible routes for the stimulated Raman transition to happen via any intermediate state. In that context, the effective Rabi frequency of the stimulated Raman transition between the hyperfine ground-states and is obtained by summing the contributions of all possible transitions involving different intermediate states. Whichever transitions are allowed depends on the polarization of the Raman lasers. Assuming both beams have the same polarisation, this leads to with the magnetic quantum number in the ground-state Weiss et al. (1994). This, with (most interferometers operate on magnetic-insensitive states), is the expression we have used in this work.